8. Select the choice below that could be the first step in solving the following system of equations using elimination: -x + 4y = 9 3x - 5y = 10
A. Multiply the second equation by 4
B. Write the first equation in terms of x.
C. Replace the right side of the first equation with the second equation
D. Multiply the first equation by 3

Answer by mananth(16946)

(-x + 4y = 9) X3
-3x+12y=27
3x - 5y = 10


D. Multiply the first equation by 3